I need the equation and step by step directions to solve this word problem.
Two cars leave a garage traveling in opposite directions. One car leaves at 8am and averages 60mph. The other car leaves at 9am and averages 50mph. At what time will they be 225 miles apart?

Answer by stanbon(75887)

Two cars leave a garage traveling in opposite directions. One car leaves at 8am and averages 60mph. The other car leaves at 9am and averages 50mph. At what time will they be 225 miles apart?
-----
Early car DATA:
rate = 60 mph ; time = x hrs ; distance = 60x miles
-----
Later car DATA:
rate = 50 mph ; time = x-1 hrs ; distance = 50(x-1) miles
-----
Equation:
distance + distance = 225 miles
60x + 50(x-1) = 225
110x = 275
x = 2.5 hrs
----
Ans: 8:00 AM + 2.5 hrs = 10:30 AM
===========
Cheers,
Stan H.